# Evalet: Evaluating Large Language Models by Fragmenting Outputs into Functions

### v2

@@ Abstract.
This figure illustrates how Evalet evaluates LLM outputs using functional fragmentation. An example output on the left is broken into highlighted text fragments, which are linked to evaluation criteria such as Engagement and Age Appropriateness. The system identifies each fragment’s function, for example, Engagement via auditory cues like “they look around and shout,” and, from Age Appropriateness perspective, introducing war-related imagery through phrases like “shooting their special microscopic guns.” Evalet rates whether each function satisfies the criterion. Functions from multiple outputs are embedded into a visualization space, enabling interpretation and validation at scale.
@@ 1. Introduction
Large Language Models (LLMs) have enabled practitioners (e.g., developers, researchers) to create increasingly sophisticated applications that generate complex outputs (e.g., stories (Chung et al., 2022; Yuan et al., 2022), research papers (Lu et al., 2024; Starace et al., 2025), and reasoning traces (Jaech et al., 2024; Snell et al., 2024)). Deploying these models safely requires rigorous verification that the outputs align (Shen et al., 2024) with practitioners’ intended goals. Evaluation is frequently manual as the applications are novel—lacking established benchmarks—and involve subjective aspects that require qualitative judgments (Kim et al., 2024a), like how insightful or harmful the application’s outputs are. Identifying systemic and recurring issues requires reviewing hundreds of outputs, but the burden of manual inspection often leads practitioners to overgeneralize from small samples (Kim et al., 2024a; Arawjo et al., 2024; Szymanski et al., 2024; Shankar et al., 2024). To address this, practitioners have begun employing LLM-based evaluators (i.e., LLM-as-a-Judge (Zheng et al., 2023)), where one LLM evaluates another’s outputs. By describing multiple criteria (e.g., Insightfulness, Harmlessness) in natural language, practitioners can assess the alignment of the model outputs with their various goals (Kim et al., 2023b; Zhong et al., 2022; Fu et al., 2023; Liu et al., 2023a).
@@
Current LLM-as-a-Judge approaches use holistic scores, where an entire output is summarized into numeric ratings (e.g., 3 out of 5) for each criterion. Holistic scores help practitioners quickly assess overall performance (Kim et al., 2024a) but obscure the specific elements in the outputs that led to these assessments. For example, in Figure 1, an LLM explaining “T cells” to a young child received a moderate score for the criterion Age Appropriateness. To understand this rating, users must manually review the output to notice that while it uses simple vocabulary, it also employs potentially harmful war imagery. This manual process provides necessary insights but undermines the automation benefits. While some LLM evaluators provide brief justifications, practitioners must still map the justification to the specific fragments in the output (Kim et al., 2024a). This lack of detail or granularity becomes more critical at scale. When multiple outputs receive identical scores, practitioners have to read the justifications for each output’s evaluation to determine whether they share the same issues or different ones. Ultimately, the opaqueness of holistic scores inhibits practitioners from identifying systemic failure patterns in the outputs that require urgent attention (Cabrera et al., 2023b; Ribeiro et al., 2020), and validating the accuracy and consistency of the LLM evaluator’s judgments (Gebreegziabher et al., 2025).
@@
To address these challenges, we propose functional fragmentation (Fig. 1): a novel LLM-based evaluation method that dissects each output into key fragments and interprets the functions of each fragment, where each fragment may serve multiple functions. With functions, we refer to the rhetorical roles or purposes that text fragments serve that are relevant to a given evaluation criterion. In Figure 1, the fragment describing T cells “shooting their special microscopic guns” serves the function of “personification” for the criterion Engagement, but also “war-related imagery” for Age Appropriateness.
@@ 2. Related Work
This work aims to support interactive evaluation of LLMs through sensemaking of model outputs and evaluations at scale. To this end, we review literature in (1) interactive evaluation for machine learning (ML), (2) interactive testing and evaluation of LLMs, and (3) sensemaking of LLM outputs at scale.
2.1. Interactive Machine Learning Evaluation
Traditionally, Machine Learning (ML) models are evaluated on benchmarks with automated metrics, where performance is aggregated into a single statistic or score. However, this provides limited signal into how the model behaves, what its flaws are, and what are the specific areas for improvement (Ribeiro et al., 2020; Zeng et al., 2025; Murahari et al., 2023). Instead of relying on aggregated metrics, prior work has introduced various systems that support interactive evaluation of models at a more fine-grained level. For example, Polyjuice (Wu et al., 2021) and AdaTest (Ribeiro and Lundberg, 2022) allow practitioners to iteratively evaluate models by creating challenging input data and testing how the models behave on these cases. Furthermore, researchers have proposed various tools (Cabrera et al., 2023a; Wexler et al., 2020; Wu et al., 2019; Robertson et al., 2023; Sivaraman et al., 2025) that help practitioners to unpack evaluations by identifying slices or subsets of data, and testing models on these to identify specific flaws or limitations. This significant body of work highlights the importance and need for more granular analysis of model performance. Our work extends this to the evaluation of LLMs: instead of simply slicing datasets, we slice the data points themselves into fragment-level functions to provide a more fine-grained understanding of LLM outputs.
@@ 2.2. Interactive Testing and Evaluation of LLMs
The general-purpose capabilities of LLMs have enabled novel AI-based applications, but have also increased the difficulty in verifying that these models perform as intended. Specifically, as these models are applied to new tasks and contexts, there are no benchmarks or metrics to automate evaluation (Kim et al., 2024a) and, as their input and output space is near infinite, the models have to be tested with numerous and diverse samples (Zamfirescu-Pereira et al., 2023; Liu et al., 2023b). To help practitioners test and evaluate these models, researchers have proposed novel tools. For example, AI Chains (Wu et al., 2022) allows users to decompose complex tasks into a chain of fine-grained inputs that can be tested individually, and ChainForge (Arawjo et al., 2024; Zhang and Arawjo, 2024) expands on this by allowing users to construct complex pipelines to test hypotheses and analyze results with various LLMs and inputs. Another body of work (Strobelt et al., 2023; Wu et al., 2023b; Mishra et al., 2023; Kim et al., 2023a) supports users to create variations of model inputs to test and compare model performance on these inputs. More recently, the success of LLM-as-a-Judge (Zheng et al., 2023) (i.e., LLMs evaluating other LLMs) has led to several systems (Kim et al., 2024a; Kahng et al., 2024; Ashktorab et al., 2024; Shankar et al., 2024) that employ LLM-based evaluators to support interactive evaluation on diverse aspects or criteria. By assessing outputs on multiple criteria, these approaches offer a multi-dimensional view of model performance. However, as they only provide holistic scores and overall justifications, practitioners must manually review the outputs and justifications to identify specific strengths and weaknesses and validate evaluations (Kim et al., 2024a; Gebreegziabher et al., 2025)—requiring effort that is impractical at scale. To address this, our work proposes employing LLMs to evaluate the individual fragments—and their functions—of outputs and grouping these evaluations across multiple outputs to surface common strengths and weaknesses, and facilitate verification of evaluations’ consistency.
2.3. Sensemaking of LLM Outputs at Scale
@@
Substantial work has explored how to support large-scale sensemaking of text artifacts to identify patterns and distill insights (Lam et al., 2024b; Wang et al., 2023; Glassman et al., 2015; Yatani et al., 2011; Kim et al., 2016; Palani et al., 2021; Rachatasumrit et al., 2021; Palani et al., 2022). Recent work has started to investigate how to support sensemaking over LLM outputs (Suh et al., 2023; Jiang et al., 2023) to facilitate exploration and analysis of diverse outputs. For example, Luminate (Suh et al., 2024) guides LLMs to generate outputs according to key dimensions and then visualizes these outputs according to these dimensions, helping writers explore the space of plausible outputs. Gero et al. (Gero et al., 2024) explored various designs and algorithms (e.g., unique words, exact matches) to support comparison of LLM outputs and help users form mental models of how the LLM behaves. Most similar to our work, Policy Projector (Lam et al., 2024a) “maps” LLM input-output pairs into a 2D space to help users explore common groups of outputs, classify these groups, and define policies or rules on the model’s behaviors based on these classified groups. In our work, we build on these approaches to propose a novel approach for multi-dimensional and granular sensemaking of LLM outputs. Instead of visualizing entire outputs, we extract fragment-level functions from multiple outputs relevant to each criterion, and then individually visualize the space of functions for each criterion—allowing users to explore fine-grained model behaviors within dimensions of interest.
@@ 3. Functional Fragmentation: An Evaluation Approach
To evaluate the alignment of an LLM, practitioners must not only quantify quality through numeric scores, but also qualitatively understand how the LLM is composing outputs and their various characteristics (Ribeiro et al., 2020; Dunlap et al., 2024; Gero et al., 2024). While existing approaches for LLM-based evaluations (Zheng et al., 2023; Kim et al., 2024a; Ye et al., 2023) can assess outputs on various dimensions or criteria, they assess outputs holistically by providing overall scores and justifications. As a result, practitioners must still inspect the outputs to further understand the specific elements of each output, and how this may satisfy or violate their goals.
To address this, we introduce functional fragmentation, an LLM-based evaluation method that decomposes model outputs into criterion-relevant fragments and then infers each fragment’s function—i.e., the role or effect it serves that influences the output’s fulfillment of that criterion. Our approach draws inspiration from inductive coding (Thomas, 2006) (i.e., interpreting raw data into codes aligned with higher-level themes) and rubric design (Perlman, 2003) (i.e., inspecting artifacts to define quality aspects to review). This section details the affordances functional fragmentation enables for inspection, In this section, we outline the novel affordances that are supported by this approach for inspecting, rating, and comparison of LLM outputs.
@@ 3.1. Inspect
To qualitatively understand the fragments present in LLM outputs, existing LLM-based evaluation approaches require practitioners to manually review the outputs and evaluator’s justifications, create mappings between these, and interpret each fragment’s significance in terms of the evaluation criteria (Kim et al., 2024a). Our approach automatically disentangles LLM outputs into key fragments relevant to a criterion and automatically interprets their functions with respect to this criterion, directly presenting practitioners with qualitative interpretations that they can inspect and verify. Additionally, as the same fragment can serve multiple functions under different criteria, our approach allows practitioners to examine the same content from multiple perspectives to gain deeper insights and even identify trade-offs. Given that criteria are often subjective, our approach can also uncover meaningful functions that the practitioner may have not previously considered—similar to the process of inductive coding (Thomas, 2006). Conversely, if the LLM evaluator extracts functions that the practitioner considers irrelevant to the criterion, practitioners can directly flag these to be ignored in future evaluations.
@@
Beyond identifying each fragment-level function in an output, practitioners may also need to inspect how these functions appear together in the output. For example, when evaluating Tension in LLM-generated horror stories, a practitioner may need to understand how the LLM uses various functions to gradually build tension in the story. Traditionally, this would require the practitioner to read the whole story, but not all of the content may be directly related to that criterion. With functional fragmentation, each output can be summarized into a list of functions related to a given criterion, allowing practitioners to easily inspect each output by focusing only on the aspects of interest.
@@ 3.2. Rate
By disentangling outputs into fragment-level functions, each individual function’s alignment with a criterion can be rated independently. More fine-grained evaluations can support interpretability by clearly highlighting the specific aspects of an output that are aligned or misaligned with a criterion. Instead of correcting the LLM evaluator by editing criteria descriptions, practitioners can directly re-rate specific functions to control future evaluations—similar to how educators develop rubrics by assessing examples of student work (Perlman, 2003). Beyond affordances for practitioners, LLMs evaluators have been shown to yield more consistent results when performing more fine-grained evaluations through checklists (Saad-Falcon et al., 2024; Lin et al., 2024) or rubrics (Kim et al., 2023b; Ye et al., 2023; Kim et al., 2024b). However, while these approaches rely on pre-defined fine-grained elements, fragment-level functions in our approach are emergent, identified dynamically based on the output and criterion.
@@ 3.3. Compare
While comparing fragments across outputs can help practitioners uncover prevalent model behaviors, comparing raw fragments is challenging. Specifically, fragments from each output may contain different lexical or semantic content, even if they serve similar functions in their respective outputs. For instance, we are evaluating an essay-writing LLM on Logical Coherence and identify these two sentences: “In conclusion, the data shows a clear declining trend in birth rate.”, and “To sum up, we should invest in public infrastructure.” Despite wording differences, in the scope of Logical Coherence, both sentences function as cohesive devices for a conclusion. By labeling each fragment’s functions, our approach allows for comparison and grouping of fragments not based on their lexical similarity, but by their functional similarity—allowing practitioners to distill high-level insights and patterns.
@@
This figure shows Evalet’s main interface, which consists of two coordinated views: an Information Panel on the left and a Map Visualization on the right. The Explore Tab in the Information Panel displays evaluation results grouped by semantic clusters for a selected criterion. Users can switch between tabs through the Tab Navigator, initiate evaluation by the Run Evaluation button, and adjust what is shown on the map via Map Controls. The Map Visualization presents each function as a point (positive) or cross (negative), clustered by semantic similarity. Hovering over a point reveals its function.
@@ 4. Evalet: Evaluation of LLM Outputs based on Fragment-Level Functions
Fragment-Level Functions: Evalet extracts all fragments relevant to a criterion and, by interpreting the role or effect that they perform, provides each fragment with a short label that describes the function it represents. Each function is given a rating (i.e., “positive” or “negative”) based on its alignment with the criterion222A single fragment can be interpreted to serve different functions for different criteria, where one such function aligns with its respective criterion while the other misaligns with its criterion. As a result, we rate each function rather than each fragment..

### v3

@@ 1. Introduction
Large Language Models (LLMs) have enabled practitioners (e.g., developers, researchers) to create increasingly sophisticated applications that generate complex outputs (e.g., stories (Chung et al., 2022; Yuan et al., 2022), research papers (Lu et al., 2024; Starace et al., 2025), and reasoning traces (Jaech et al., 2024; Snell et al., 2024)). Deploying these models safely requires rigorous verification that the outputs align (Shen et al., 2024) with practitioners’ intended goals. Evaluation is frequently manual as the applications are novel—lacking established benchmarks—and involve subjective aspects that require qualitative judgments (Kim et al., 2024b), like how insightful or harmful the application’s outputs are. Identifying systemic and recurring issues requires reviewing hundreds of outputs, but the burden of manual inspection often leads practitioners to overgeneralize from small samples (Kim et al., 2024b; Arawjo et al., 2024; Szymanski et al., 2024; Shankar et al., 2024). To address this, practitioners have begun employing LLM-based evaluators (i.e., LLM-as-a-Judge (Zheng et al., 2023)), where one LLM evaluates another’s outputs. By describing multiple criteria (e.g., , ) in natural language, practitioners can assess the alignment of the model outputs with their various goals (Kim et al., 2023a; Zhong et al., 2022; Fu et al., 2023; Liu et al., 2023b).
@@
Current LLM-as-a-Judge approaches use holistic scores, where an entire output is summarized into numeric ratings (e.g., 3 out of 5) for each criterion. Holistic scores help practitioners quickly assess overall performance (Kim et al., 2024b) but obscure the specific elements in the outputs that led to these assessments. For example, in Figure 1, an LLM explaining “T cells” to a young child received a moderate score for the criterion . To understand this rating, users must manually review the output to notice that while it uses simple vocabulary, it also employs potentially harmful war imagery. This manual process provides necessary insights but undermines the automation benefits. While some LLM evaluators provide brief justifications, practitioners must still map the justification to the specific fragments in the output (Kim et al., 2024b). This lack of detail or granularity becomes more critical at scale. When multiple outputs receive identical scores, practitioners have to read the justifications for each output’s evaluation to determine whether they share the same issues or different ones. Ultimately, the opaqueness of holistic scores inhibits practitioners from identifying systemic failure patterns in the outputs that require urgent attention (Cabrera et al., 2023b; Ribeiro et al., 2020), and validating the accuracy and consistency of the LLM evaluator’s judgments (Gebreegziabher et al., 2025).
@@
To address these challenges, we propose functional fragmentation (Fig. 1): a novel LLM-based evaluation method that dissects each output into key fragments and interprets the functions of each fragment, where each fragment may serve multiple functions. With functions, we refer to the rhetorical roles or purposes that text fragments serve that are relevant to a given evaluation criterion. In Figure 1, the fragment describing T cells as “shooting their special microscopic guns” serves the function of “personification” for the criterion , but also “war-related imagery” for .
@@ 2. Related Work
This work aims to support interactive evaluation of LLMs through sensemaking of model outputs and evaluations at scale. To this end, we review literature in (1) interactive examination of machine learning (ML) models, (2) interactive testing and evaluation of LLMs, and (3) sensemaking of text at scale.
2.1. Interactive Examination of Machine Learning
Traditionally, Machine Learning (ML) models are examined and understood by evaluating on benchmarks with automated metrics that aggregate performance into a single statistic or score. However, this provides limited understanding into how the model behaves, what its flaws are, and what are the areas for improvement (Ribeiro et al., 2020; Zeng et al., 2025; Murahari et al., 2023). Instead of relying on aggregated metrics, prior work has introduced systems that support more interactive and fine-grained evaluation of models. For example, Polyjuice (Wu et al., 2021) and AdaTest (Ribeiro and Lundberg, 2022) allow practitioners to iteratively evaluate models by creating challenging input data and testing how the models behave on these cases. Furthermore, researchers have proposed various tools (Cabrera et al., 2023a; Wexler et al., 2020; Wu et al., 2019; Robertson et al., 2023; Sivaraman et al., 2025) that help practitioners to unpack evaluations by identifying slices or subsets of data, and testing models on these to identify specific flaws or limitations. Complementing these evaluation approaches, a rich body of work in explainable AI (XAI) (Liao and Varshney, 2021) has also explored how to support understanding of models through more fine-grained analysis of behaviors (Lai et al., 2022). For instance, the foundational methods LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017) explain model predictions by inferring the effect of individual features. Frameworks such as the XAI Question Bank (Liao et al., 2020) organizes users’ understanding and explainability needs into fine-grained questions that explore diverse aspects of models (e.g., inputs, outputs, performance). Collectively, this work highlights the importance and need for more granular analysis of model performance. Our work extends this to the evaluation of LLMs: instead of simply slicing datasets, we slice the data points themselves into fragment-level functions to provide a more fine-grained understanding of LLM outputs.
@@ 2.2. Interactive Testing and Evaluation of LLMs
The general-purpose capabilities of LLMs have enabled novel AI applications but also made it harder to verify that they behave as intended. As these models are applied to new tasks and contexts, there are no benchmarks or metrics to automate evaluation (Kim et al., 2024b) and, due to their near infinite input-output space, models have to be tested with numerous and diverse samples (Zamfirescu-Pereira et al., 2023; Liu et al., 2023a). To help practitioners, researchers have proposed novel tools that support interactive testing and experimentation on these models by decomposing tasks into chains of sub-tasks (Wu et al., 2022), composing diverse LLM pipelines in parallel (Arawjo et al., 2024; Zhang and Arawjo, 2024), or creating diverse variations of test inputs (Strobelt et al., 2023; Wu et al., 2023a; Mishra et al., 2023; Kim et al., 2023b). More recently, the success of LLM-as-a-Judge (Zheng et al., 2023) (i.e., LLMs evaluating other LLMs) has led to several systems (Kim et al., 2024b; Kahng et al., 2024; Ashktorab et al., 2024; Shankar et al., 2024) that employ LLM-based evaluators to support interactive evaluation on diverse criteria—offering a multi-dimensional view of model performance. However, as these only provide holistic scores and overall justifications, practitioners must manually review the outputs and justifications to identify specific strengths and weaknesses, and validate the evaluations (Kim et al., 2024b; Gebreegziabher et al., 2025)—requiring effort that is impractical at scale. To address the limitations of holistic scoring, prior work has proposed more fine-grained evaluation methods. For example, Nenkova and Passonneau (Nenkova and Passonneau, 2004) and FactScore (Min et al., 2023) decompose text into units, which are then assessed individually, while Scarecrow (Dou et al., 2021) and BooookScore (Chang et al., 2023) assess outputs by annotating spans on predefined error categories (Dou et al., 2021; Chang et al., 2023). Building on this, our approach employs LLMs to decompose outputs into fragments, annotate their function in an emergent manner, and cluster these across outputs—surfacing common strengths and weaknesses, and facilitating verification of evaluation consistency.
2.3. Sensemaking of Text at Scale
@@
Theories on sensemaking posit that people make sense of large information spaces through multiple cognitive processes such as iteratively foraging and organizing information—the notional model (Pirolli and Card, 2005)—and comparing information to identify patterns—structure-mapping theory (Gentner, 1983). Given their significant cognitive demands, prior work has proposed systems that support these processes for textual data: facilitating structuring and organization of collected information (Palani et al., 2022; Rachatasumrit et al., 2021), generating summaries or topic models (Lam et al., 2024b; Yatani et al., 2011; Palani et al., 2021; Glassman et al., 2015), and visualizing corpora using spatial embeddings or structural patterns (Kim et al., 2016; Wang et al., 2023; Gu et al., 2025). Recent work extends these ideas to sensemaking over LLM outputs (Suh et al., 2023; Jiang et al., 2023). For example, Luminate (Suh et al., 2024) guides LLMs to generate outputs along key dimensions and then visualizes the outputs on these dimensions, helping writers explore the generation space. Gero et al. (Gero et al., 2024) explored various designs and algorithms (e.g., unique words, exact matches) to support comparison of LLM outputs and help users form mental models of LLM behavior. Most similar to our work, Policy Projector (Lam et al., 2024a) “maps” LLM input-output pairs into a 2D space to help users explore common groups of outputs, classify these groups, and define policies on the model’s behaviors using these classified groups. Building on these approaches, we propose a novel approach for multi-dimensional and granular sensemaking of LLM outputs. Instead of visualizing entire outputs, we extract fragment-level functions from multiple outputs for each criterion and then visualize the space of functions for each criterion—supporting exploration of fine-grained model behaviors within dimensions of interest.
@@ 3. Functional Fragmentation: An Evaluation Approach
To evaluate the alignment of an LLM, practitioners must not only quantify quality through numeric scores but also qualitatively understand how this models’ outputs are composed and their characteristics (Ribeiro et al., 2020; Dunlap et al., 2024; Gero et al., 2024). While existing LLM-based evaluation methods can assess outputs across various criteria (Zheng et al., 2023; Kim et al., 2024b; Ye et al., 2023), they only provide holistic judgments (i.e., overall scores, justification) for each dimension. Thus, practitioners must manually inspect outputs to identify the specific elements that satisfy or violate their goals.
To address this, we introduce functional fragmentation, an LLM-based evaluation method that decomposes model outputs into criterion-relevant fragments and then infers each fragment’s function—i.e., the role or effect it serves that influences the output’s fulfillment of that criterion. Our approach draws inspiration from inductive coding (Thomas, 2006) (i.e., interpreting raw data into codes based on higher-level themes) and rubric design (Perlman, 2003) (i.e., inspecting artifacts to define quality aspects to review). In this section, we outline the novel affordances that are supported by this approach for inspecting, rating, and comparison of LLM outputs.
@@ 3.1. Inspect
To make sense of existing LLM-based evaluations, practitioners must manually review outputs, connect them with the evaluator’s justifications, and interpret each fragment’s significance in terms of the criteria (Kim et al., 2024b). Our approach automatically extracts criteria-relevant fragments and interprets their functions with respect to each criterion—directly presenting practitioners with qualitative interpretations to inspect and verify. As one fragment can serve different functions under different criteria, our approach allows practitioners to examine the same content from multiple perspectives and identify trade-offs. Given that criteria are often subjective, our approach can also uncover meaningful functions that the practitioner may have not previously considered—similar to inductive coding (Thomas, 2006). Conversely, if the LLM evaluator extracts functions that the practitioner considers irrelevant to the criterion, practitioners can directly flag these to be ignored in future evaluations.
@@
Beyond identifying each fragment-level function in an output, practitioners may also need to inspect how these functions appear together in the output. For example, when evaluating in LLM-generated horror stories, a practitioner may need to understand how the LLM uses various functions to gradually build tension in the story. Traditionally, this would require the practitioner to read the whole story, but not all of the content may be directly related to that criterion. With functional fragmentation, each output can be summarized into a list of functions related to a given criterion, allowing practitioners to easily inspect each output by focusing only on the aspects of interest.
@@ 3.2. Rate
By disentangling outputs into fragment-level functions, each individual function’s alignment with a criterion can be rated independently. More fine-grained evaluations can support interpretability by clearly highlighting the specific aspects of an output that are aligned or misaligned with a criterion. Instead of correcting the LLM evaluator by editing criteria descriptions, practitioners can directly re-rate specific functions to control future evaluations—similar to how educators develop rubrics by assessing examples of student work (Perlman, 2003). Beyond supporting practitioners, LLM evaluators are also more consistent when performing more fine-grained evaluations through checklists (Saad-Falcon et al., 2024; Lin et al., 2024) or rubrics (Kim et al., 2023a; Ye et al., 2023; Kim et al., 2024a). Unlike these approaches, which rely on predefining these checklists and rubrics, our fragment-level functions are emergent, identified dynamically based on the outputs and criteria.
@@ 3.3. Compare
Comparing fragments across outputs can reveal common model behaviors, but directly comparing raw text is difficult because fragments may differ lexically or semantically even when they serve the similar function. For example, when we evaluate an essay-writing LLM on , these two sentences serve the same function as cohesive devices for a conclusion despite their wording differences: “In conclusion, the trend is clear.”, and “To sum up, it supports our view.” By labeling each fragment’s functions, our approach allows for comparison and grouping of fragments not based on their lexical similarity, but by their functional similarity—allowing practitioners to distill high-level insights and patterns.
@@ 4. Evalet: Evaluation of LLM Outputs based on Fragment-Level Functions
Fragment-Level Functions: Evalet extracts criterion-relevant fragments and interprets the functional role or effect of each fragment to assign a short function label. Each function receives a “positive” or “negative” rating based on its alignment with the criterion222A single fragment can be interpreted to serve different functions for different criteria, where one such function aligns with its respective criterion while the other misaligns with its criterion. As a result, we rate each function rather than each fragment..